ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 135, 2003

Alon, G. and E. de Shalit Cohomology of discrete groups in harmonic cochains on buildings

355

Argyros, S. A., A. Manoussakis and M. Petrakis Function spaces not containing ℓ_{1}

29

Ben-Itzhak, T. and M. Teicher Graph theoretic method for determining Hurwitz equivalence in the symmetric group

83

Berrick, A. J. and E. Dror Farjoun Fibrations and nullifications 205

Bianchi, G. and P. Gronchi Steiner symmetrals and their distance from a ball

181

Bonnet, R. See M. Rubin

de Shalit, E. See G. Alon

Downarowicz, T. and J. Serafin Possible entropy functions  221

Dror Farjoun, E. See A. J. Berrick

M. Fradelizi, A. Giannopoulos and M. Meyer Some inequalities about mixed volumes

157

Frankild, A. and P. Jørgensen Gorenstein differential graded algebras

327

Giambruno, A. and M. Zaicev Asymptotics for the standard and the Capelli identities

125

Giannopoulos, A. See M. Fradelizi

Gronchi, P. See G. Bianchi

Jones, R. L., J. M. Rosenblatt and M. Wierdl Oscillation in ergodic theory: Higher dimensional results

1

Jørgensen, P. See A. Frankild

Levin, M. Acyclic resolutions for arbitrary groups  193

Manoussakis, A. See S. A. Argyros

Medvedev, Y. On compact Engel groups  147

Meyer, M. See M. Fradelizi

Ould-Beddi, M. A. and L. Strüngmann Tight subgroups in torsion-free abelian groups

251

Petrakis, M. See S. A. Argyros Peyerimhoff, N. and K. F. Siburg The dynamics of magnetic flows

for energies above Mañé’s critical value

269

Rosenblatt, J. M. See R. L. Jones Rubin, M. and R. Bonnet On a poset algebra which is hereditarily

but not canonically well generated

299

Serafin, J. See T. Downarowicz Siburg, K. F. See N. Peyerimhoff

Solecki, S. Translation invariant ideals  93

Strüngmann, L. See M. A. Ould-Beddi Teicher, M. See T. Ben-Itzhak Wierdl, M. See R. L. Jones Zaicev, M. See A. Giambruno

Zhu, C.-B. Representations with scalar K-types and applications  111